### v2

@@ 4 Swap distance and insert distance
The swap (insert) distance between two schedules $\pi$ and $\pi^{\prime}$ is then defined as the length of the shortest sequence of consecutive swap (insert) operation to transform $\pi$ into $\pi^{\prime}$.
In the following, we show that for the swap and insert distance, $\mathrm{FPT}$-algorithms for the search radius $k$ are unlikely. More precisely, we show that even if each job has processing time one and the given schedule $\pi$ tardiness zero, $\mathrm{FPT}$-algorithms for $k$ plus many other structural parameters are unlikely. On the positive side, we show that a simple brute force algorithm with running time $n^{2k+1}$ is possible for both distance measures.
Theorem 4.1
Swap LS MM and Insert LS MM are $\mathrm{W}[1]$-hard when parameterized by the search radius $k$. This holds even on instances where simultaneously
@@
We present a parameterized reduction from LocalTSP(Swap). To this end, we first introduce some notation to define LocalTSP(Swap). For a graph $G=(V,E)$, a Hamiltonian cycle is a permutation of the vertex set $C=(v_{1},\dots,v_{n})$ such that between any two consecutive vertices in $C$, there is an edge between these vertices in $G$, where we consider vertex $v_{1}$ to follow $v_{n}$. That is, the edges of $C$ defined by $E(C):=\{\{v_{i},v_{i+1}\}\mid i\in[1,n-1]\}\cup\{\{v_{n},v_{1}\}\}$ are all contained in $E$. For an edge-weight function $\omega$ and a Hamiltonian cycle $C$, we denote by $\omega(C):=\sum_{e\in E(C)}\omega(e)$ the total weight of all edges of $C$. LocalTSP(Swap) is now formally defined as follows.
@@
We now define the initial schedule $\pi$ as $\pi:=\pi_{C}$ and $k^{\prime}:=4k$. This completes the construction of $I^{\prime}$. The intuitive idea is that each schedule that improved over $\pi$ is equal to $\pi_{C^{\prime}}$ for some Hamiltonian cycle $C^{\prime}$ for $G$ which improves over $C$ and vice versa. Moreover, the swap-distance between $\pi=\pi_{C}$ and $\pi_{C^{\prime}}$ is at most four times the swap-distance between $C$ and $C^{\prime}$.
@@
Since there are four jobs for each vertex of $V$, the swap-distance between $\pi$ and $\pi_{C^{\prime}}$ is at most four times the swap-distance between $C$ and $C^{\prime}$. $\blacksquare$
@@
Due to Section 4, this is at least the total setup time of $\pi$ and thus, $\pi^{\prime}$ does not improve over $\pi$. $\blacksquare$
@@
Proof. Since $\pi^{\prime}$ improves over $\pi$, Section 4 implies that $\pi^{\prime}$ is light. Moreover, Section 4 further implies that there is a vertex $u\in V$ such that $\pi^{\prime}$ starts with $v_{1}^{\mathrm{r}}$ ($v_{1}^{\mathrm{l}}$) followed by $v_{1}^{\mathrm{out}}$ ($v_{1}^{\mathrm{in}}$) and ends with $v_{1}^{\mathrm{l}}$ ($v_{1}^{\mathrm{r}}$) preceded by $v_{1}^{\mathrm{in}}$ ($v_{1}^{\mathrm{out}}$). Recall that the setup matrix is symmetrical and that all jobs have a deadline of $\infty$. Hence, a schedule $\pi^{\prime}$ and its reverse schedule both have the same makespan and the same tardiness of zero. Without loss of generality, we may thus assume that $\pi^{\prime}$ starts with $v_{1}^{\mathrm{r}}$ followed by $v_{1}^{\mathrm{out}}$ and ends with $v_{1}^{\mathrm{l}}$ preceded by $v_{1}^{\mathrm{in}}$.
@@
Note that this implies that any schedule $\pi^{\prime}$ for $J$ has total setup time at least $39n-$ 3 due to Section 4. Hence, the makespan of $\pi$ can be improved by at most 1. Moreover, this implies that each schedule that improves over $\pi$ is a globally optimal schedule.
@@
()($\Rightarrow$)( ⇒ ) Let $C^{\prime}$ be a Hamiltonian cycle for $G$ with $\omega(C^{\prime})<\omega(C)$ and where $C$ and $C^{\prime}$ are $k$-swap-neighbors. Since $\omega(C)=1$, this implies that $\omega(C^{\prime})=0$. Consider the schedule $\pi_{C^{\prime}}$ of $J$. Due to Section 4, the total setup time of $\pi_{C^{\prime}}$ is $39n-$ 3 and $\pi$ and $\pi_{C^{\prime}}$ are $k^{\prime}$-swap-neighbors. Hence, $\pi_{C^{\prime}}$ improves over $\pi$ and thus, $I^{\prime}$ is a yes-instance of Swap LS MM, where a globally optimal solution is a $k^{\prime}$-swap-neighbor of the initial schedule.
()($\Leftarrow$)( ⇐ ) Let $\pi^{\prime}$ be a schedule of $J$ that improves over $\pi$ where $\pi$ and $\pi^{\prime}$ are $k^{\prime}$-swap-neighbors. Due to Section 4, we can assume without loss of generality, that $\pi^{\prime}$ is equal to $\pi_{C^{\prime}}$ for some Hamiltonian cycle $C^{\prime}$ of $G$. Since $\pi^{\prime}$ improves over $\pi$, Section 4 implies that $\omega(C^{\prime})<\omega(C)$. Hence, $C$ is not a globally optimal solution for $I$ which implies that $C$ is not $k$-swap-optimal, since the instance $I$ of LocalTSP(Swap) provides the property that $C$ is a globally optimal solution if and only if $C$ is $k$-swap-optimal. Consequently, $I$ is a yes-instance of LocalTSP(Swap). $\square$
In other words, even on very restricted instances and for each computable function $f$, a running time of $f(k)\cdot|I|^{\mathcal{O}(1)}$ for Swap LS MM or Insert LS MM is unlikely. Still, a simple brute-force algorithm yields the following running time for both problems.
Theorem 4.2
@@
5 Derived hill climbing algorithms
The Win and MW Algorithms.
The general idea of these hill climbing algorithms is to start with some initial solution $\pi$ and $k=4$. The algorithm then tries to replace the current solution by a better one with window distance or multi-window distance at most $k$. To evaluate whether there is a better solution with distance at most $k$, we use the algorithm behind Theorems 3.1 and 3.2. If no such better solution exists, the algorithm is stuck in a locally optimal solution. Aiming to escape this locally optimal solution, the value of $k$ is incremented, which yields a larger neighborhood to consider. Otherwise, if there is a better schedule $\pi^{\prime}$ with distance at most $k$, we replace $\pi$ by $\pi^{\prime}$ and set $k$ back to $4$. We always set $k$ back to $4$ after finding any better solution, because searching for a better solution with distance at most $k$ can be performed faster the smaller $k$ is. We chose the reset value of $k$ to be $4$ instead of $2$ because preliminary experiments showed that for $k\in\{2,3\}$, improvements happened rarely.
The Win+Swap and MW+Swap Algorithms.
The algorithms Win+Swap and MW+Swap extend Win and MW by additionally searching for better solutions with swap distance 1. We limit the algorithms to only search for swap distance 1, since the hardness result of Theorem 4.1 implies that considering larger swap distances tends to become inefficient. This was also validated by preliminary experiments when considering also better solutions of swap distance 2 and 3. Further, preliminary experiments showed that considering the swap distance provided much better solutions than considering the insert distance. As a consequence, we do not include combinations of insert distance with any of window distance or multi-window distance in our evaluation.
The pseudo code for Win+Swap is depicted in Algorithm 1. Lines 47 of Algorithm 1 check for a better schedule in the $1$-swap-neighborhood. ,color=yellow!90!red]Diese Erklärungen brauchen wir doch jetzt nicht mehr, oder? Omitting these lines yields our second algorithm: Win. By replacing in Line 8 and Line 11 the window distance by the multi-window distance yields MW+Swap. The fourth local search algorithm (MW) is obtained by performing both these operations, that is, omitting Lines 47 and replacing in Line 8 and Line 11 the window distance by the multi-window distance.
Tardiness and starting solutions.
We next describe the starting solutions we used for our algorithms. Unfortunately, providing a feasible starting solution can not be done in polynomial time, unless $\mathrm{P}=\mathrm{NP}$. Hence, in our experimental evaluation, we also allow non-feasible schedules. Then, our goal is to minimize the total tardiness and the makespan, where minimizing the total tardiness is the primary objective. That is, a schedule $\pi^{\prime}$ is better than a schedule $\pi$, if a) the total tardiness of $\pi^{\prime}$ is smaller than the total tardiness of $\pi$ or b) $\pi$ and $\pi^{\prime}$ have the same total tardiness and the makespan of $\pi^{\prime}$ is smaller than the makespan of $\pi$. Note that under this modified objective function, if the current solution $\pi$ has total tardiness zero, the algorithms for all distance measures are still correct. If the total tardiness of $\pi$ is non-zero, the previous described algorithms for finding better solution with window distance or multi-window distance are not necessarily correct anymore, that is, the described algorithms output that there is no better solution with distance at most $k$ even though there might be a better solution with distance at most $k$, where the corresponding change is not EDDS. In our evaluation, we show that even though some local improvements might not be found, the approaches still perform very well when additionally aiming to minimize the total tardiness also for non-feasible solutions.
@@ 6.1 Baseline Algorithms
More precisely, until a given time limit is reached, the algorithm takes a schedule $\pi$ from $\mathrm{ND}$ uniformly at random and tries to find a schedule $\pi^{\prime}$ which is not dominated by any schedule of $\mathrm{ND}$ such that $\pi$ and $\pi^{\prime}$ have swap-distance 1 or insert-distance 1. If such a solution $\pi^{\prime}$ is found, $\pi^{\prime}$ is added to $\mathrm{ND}$ and all schedules of $\mathrm{ND}$ that are dominated by $\pi^{\prime}$ are removed from $\mathrm{ND}$.
@@
For a detailed introduction into genetic algorithm, we refer to the standard monograph [14]. As a baseline, we additionally consider two genetic algorithms which maintain a population of size $p$ and consist of standard operators Encode ($E_{n}$), Decode ($D_{e}$), Crossover ($C_{o}$), Mutation ($M_{u}$), Evaluate ($E_{v}$), and Selection ($S_{e}$). We propose a basic genetic algorithm (GAD) and a modified genetic algorithm (MGA). To evaluate the quality of the results, we define the fitness function as $\Phi(\pi):=C_{\max}+\mu T$, where $T$ is the total tardiness of $\pi$, with $\mu:=n\cdot(\max_{j\in J}t_{j}+\max_{i,j\in J}\ell(\tau_{i},\tau_{j}))+1$ for the job set $J$. Note that $\Phi(\pi)$ essentially models a lexicographic optimization, where the primary goal is to minimize the total tardiness and the secondary goal is to minimize the setup time.
@@ 6.2 Experiments
We perform experiments to evaluate our approaches for the real-world use case from the company CUNA. The products produced by CUNA come in eight different color/material combinations, all of which are manufactured on the same injection molding machine. For our experiments, we use the $8\times 8$ setup matrix containing the machines setup times in minutes as integers.
@@
For all three starting solutions (DD, SM, TM), we evaluated our four hill climbing strategies (Win, Win+Sw, MW, MW+Sw) on 1)1)1 ) a dataset of instances where feasible schedules exist, and 2)2)2 ) a dataset of instances where no feasible schedule exists. Each instance ran within a time limit of 20 minutes. Furthermore, as a baseline, we ran experiments with the two genetic algorithms GAD and MGA and with the PILS1 algorithm. Since GAD, MGA, and PILS1 are non-deterministic randomized algorithms, we performed 20 independent repetitions with a time limit of 20 minutes each and kept the best found results.
We summarize our main experimental findings as follows: First, all hill climbing algorithms provide good results for instances where feasible schedules are possible. The algorithms were mostly able to provide good results even if the starting solution was not feasible. Overall, the window algorithm appears to be well-suited to handle these instances. Second, on instances where no feasible schedules exist, local search via single swaps often has the biggest influence on the solution quality and provides good results. Especially for the large instances with $n\in\{250,300\}$, all improvements were reached by swaps and no parameterized local search step has been applied within the time limit of 20 minutes. On the smaller instances, the VNS algorithms Win+Sw and MW+Sw help the local search via swaps to escape bad swap-optimal solutions and provide good results for this case. Overall, Win+Sw appears to be well-suited to handle these instances. Third, the hill climbing algorithms outperform the baseline genetic algorithms in the solution quality. On the small instances ($n=50$), PILS1 performed slightly better than the hill-climbing algorithms, while on larger instances ($n\geq 100$), the parameterized local search algorithms mostly provide better results than PILS1. This might indicate that parameterized local search is a reasonable alternative to the perturbation of schedules to escape poor local swap-optimal solutions.

### v3

@@ 4 Swap distance and insert distance
The swap (insert) distance between two schedules $\pi$ and $\pi^{\prime}$ equals $|\mathcal{S}|-1$, where $\mathcal{S}$ is a shortest sequence of schedules $(\pi,\dots,\pi^{\prime})$ that consecutively have swap (insert) distance one.
In the following, we show that for the swap and insert distance, $\mathrm{FPT}$-algorithms for the search radius $k$ are unlikely. More precisely, we show that even if each job has processing time one and the given schedule $\pi$ tardiness zero, $\mathrm{FPT}$-algorithms for $k$ plus further natural structural parameters are unlikely. On the positive side, we show that a simple brute force algorithm with running time $n^{2k+1}$ is possible for both distance measures.
Theorem 5.
Unless $\mathrm{FPT}=\mathrm{W}[1]$, neither of Swap LS MM or Insert LS MM admits an $\mathrm{FPT}$-algorithm when parameterized by the search radius $k$. This holds even on instances where simultaneously
@@
We present a parameterized reduction from LocalTSP(Swap) to the decision version of Swap LS MM, where the question is: Is there a better schedule with swap distance at most $k$ from the initial schedule? To this end, we first introduce some notation to define LocalTSP(Swap). For a graph $G=(V,E)$, a Hamiltonian cycle is a permutation of the vertex set $C=(v_{1},\dots,v_{n})$ such that between any two consecutive vertices in $C$, there is an edge between these vertices in $G$, where we consider vertex $v_{1}$ to follow $v_{n}$. That is, the edges of $C$ defined by $E(C):=\{\{v_{i},v_{i+1}\}\mid i\in[1,n-1]\}\cup\{\{v_{n},v_{1}\}\}$ are all contained in $E$. For an edge-weight function $\omega$ and a Hamiltonian cycle $C$, we denote by $\omega(C):=\sum_{e\in E(C)}\omega(e)$ the total weight of all edges of $C$. LocalTSP(Swap) is now formally defined as follows.
@@
We now define the initial schedule $\pi$ as $\pi:=\pi_{C}$ and $k^{\prime}:=4k$. This completes the construction of $I^{\prime}$.
The intuitive idea is that each schedule that improves over $\pi$ is equal to $\pi_{C^{\prime}}$ for some Hamiltonian cycle $C^{\prime}$ for $G$ which improves over $C$ and vice versa. Moreover, the swap distance between $\pi=\pi_{C}$ and $\pi_{C^{\prime}}$ is at most four times the swap distance between $C$ and $C^{\prime}$.
@@
Since there are four jobs for each vertex of $V$, the swap distance between $\pi$ and $\pi_{C^{\prime}}$ is at most four times the swap distance between $C$ and $C^{\prime}$. $\blacksquare$
@@
Due to 1, this is at least the total setup time of $\pi$ and thus, $\pi^{\prime}$ does not improve over $\pi$. $\blacksquare$
@@
Proof. Since $\pi^{\prime}$ improves over $\pi$, 2 implies that $\pi^{\prime}$ is light. Moreover, 3 further implies that there is a vertex $u\in V$ such that $\pi^{\prime}$ starts with $v_{1}^{\mathrm{r}}$ ($v_{1}^{\mathrm{l}}$) followed by $v_{1}^{\mathrm{out}}$ ($v_{1}^{\mathrm{in}}$) and ends with $v_{1}^{\mathrm{l}}$ ($v_{1}^{\mathrm{r}}$) preceded by $v_{1}^{\mathrm{in}}$ ($v_{1}^{\mathrm{out}}$). Recall that the setup matrix is symmetrical and that all jobs have a deadline of $\infty$. Hence, a schedule $\pi^{\prime}$ and its reverse schedule both have the same makespan and the same tardiness of zero. Without loss of generality, we may thus assume that $\pi^{\prime}$ starts with $v_{1}^{\mathrm{r}}$ followed by $v_{1}^{\mathrm{out}}$ and ends with $v_{1}^{\mathrm{l}}$ preceded by $v_{1}^{\mathrm{in}}$.
@@
Note that this implies that any schedule $\pi^{\prime}$ for $J$ has total setup time at least $39n-$ 3 due to 1. Hence, the makespan of $\pi$ can be improved by at most 1. Moreover, this implies that each schedule that improves over $\pi$ is a globally optimal schedule.
@@
()($\Rightarrow$)( ⇒ ) Let $C^{\prime}$ be a Hamiltonian cycle for $G$ with $\omega(C^{\prime})<\omega(C)$ and where $C$ and $C^{\prime}$ are $k$-swap-neighbors. Since $\omega(C)=1$, this implies that $\omega(C^{\prime})=0$. Consider the schedule $\pi_{C^{\prime}}$ of $J$. Due to 1, the total setup time of $\pi_{C^{\prime}}$ is $39n-$ 3 and $\pi$ and $\pi_{C^{\prime}}$ are $k^{\prime}$-swap-neighbors. Hence, $\pi_{C^{\prime}}$ improves over $\pi$ and thus, $I^{\prime}$ is a yes-instance of Swap LS MM, where a globally optimal solution is a $k^{\prime}$-swap-neighbor of the initial schedule.
()($\Leftarrow$)( ⇐ ) Let $\pi^{\prime}$ be a schedule of $J$ that improves over $\pi$ where $\pi$ and $\pi^{\prime}$ are $k^{\prime}$-swap-neighbors. Due to 4, we can assume without loss of generality, that $\pi^{\prime}$ is equal to $\pi_{C^{\prime}}$ for some Hamiltonian cycle $C^{\prime}$ of $G$. Since $\pi^{\prime}$ improves over $\pi$, 1 implies that $\omega(C^{\prime})<\omega(C)$. Hence, $C$ is not a globally optimal solution for $I$ which implies that $C$ is not $k$-swap-optimal, since the instance $I$ of LocalTSP(Swap) provides the property that $C$ is a globally optimal solution if and only if $C$ is $k$-swap-optimal. Consequently, $I$ is a yes-instance of LocalTSP(Swap). ∎
In other words, even on very restricted instances and for each computable function $f$, a running time of $f(k)\cdot|I|^{\mathcal{O}(1)}$ for Swap LS MM or Insert LS MM is unlikely. Still, a simple brute-force algorithm yields the following polynomial running time for each constant value of the search radius $k$ for both problems.
Theorem 6.
@@
5 Derived hill climbing algorithms
The Win and MW Algorithms.
The general idea of these hill climbing algorithms is to start with some initial solution $\pi$ and $k=4$. The algorithm then tries to replace the current solution by a better one with window distance or multi-window distance at most $k$. To evaluate whether there is a better solution with distance at most $k$, we use the algorithm behind Theorems 3 and 4. If no such better solution exists, the algorithm is stuck in a locally optimal solution. Aiming to escape this locally optimal solution, the value of $k$ is incremented, which yields a larger neighborhood to consider. Otherwise, if there is a better schedule $\pi^{\prime}$ with distance at most $k$, we replace $\pi$ by $\pi^{\prime}$ and set $k$ back to $4$. We always set $k$ back to $4$ after finding any better solution, because searching for a better solution with distance at most $k$ can be performed faster the smaller $k$ is. We chose the reset value of $k$ to be $4$ instead of $2$ because preliminary experiments showed that for $k\in\{2,3\}$, improvements happened rarely.
The Win+Swap and MW+Swap Algorithms.
The algorithms Win+Swap and MW+Swap extend Win and MW by additionally searching for better solutions with swap distance 1. We limit the algorithms to only search for swap distance 1, since the hardness result of Theorem 5 implies that considering larger swap distances tends to become inefficient. This was also validated by preliminary experiments when considering also better solutions with swap distance 2 and 3. Further, preliminary experiments showed that considering the swap distance provided much better solutions than considering the insert distance. As a consequence, we do not include combinations of insert distance with any of window distance or multi-window distance in our evaluation.
The pseudo code for Win+Swap is depicted in Algorithm 1. Lines 47 of Algorithm 1 check for a better schedule in the $1$-swap-neighborhood. ,color=yellow!90!red]Diese Erklärungen brauchen wir doch jetzt nicht mehr, oder? Omitting these lines yields our second algorithm: Win. By replacing in Line 8 and Line 11 the window distance by the multi-window distance yields MW+Swap.
@@ 5 Derived hill climbing algorithms
Tardiness.
Note that providing a feasible starting solution with tardiness zero can not be done in polynomial time, unless $\mathrm{P}=\mathrm{NP}$. Hence, in our experimental evaluation, we also allow non-feasible schedules. Then, our goal is to minimize the total tardiness and the makespan, where minimizing the total tardiness is the primary objective. That is, a schedule $\pi^{\prime}$ is better than a schedule $\pi$, if a) the total tardiness of $\pi^{\prime}$ is smaller than the total tardiness of $\pi$ or b) $\pi$ and $\pi^{\prime}$ have the same total tardiness and the makespan of $\pi^{\prime}$ is smaller than the makespan of $\pi$. Note that under this modified objective function, if the current solution $\pi$ has total tardiness zero, the algorithms for all distance measures are still correct. If the total tardiness of $\pi$ is non-zero, the previous described algorithms for finding better solution with window distance or multi-window distance are not necessarily correct anymore, that is, the described algorithms output that there is no better solution with distance at most $k$ even though there might be a better solution with distance at most $k$, where the corresponding change is not EDDS. In our evaluation, we show that even though some local improvements might not be found, the approaches still perform very well when additionally aiming to minimize the total tardiness also for non-feasible solutions.
Considered Starting Solutions.
@@ 6.1 Baseline Algorithms
More precisely, until a given time limit is reached, the algorithm takes a schedule $\pi$ from $\mathrm{ND}$ uniformly at random and tries to find a schedule $\pi^{\prime}$ which is not dominated by any schedule of $\mathrm{ND}$ such that $\pi$ and $\pi^{\prime}$ have swap distance 1 or insert distance 1 via bruteforce. If such a solution $\pi^{\prime}$ is found, $\pi^{\prime}$ is added to $\mathrm{ND}$ and all schedules of $\mathrm{ND}$ that are dominated by $\pi^{\prime}$ are removed from $\mathrm{ND}$.
@@
For a detailed introduction into genetic algorithm, we refer to the standard monograph [14]. Additionally to PILS1, we consider two genetic algorithms as a baseline. The algorithms maintain a population of size $p$ and consist of standard operators Encode ($E_{n}$), Decode ($D_{e}$), Crossover ($C_{o}$), Mutation ($M_{u}$), Evaluate ($E_{v}$), and Selection ($S_{e}$). We propose a basic genetic algorithm (GAD) and a modified genetic algorithm (MGA) that exploits the solution structure from Proposition 1. To evaluate the quality of the results, we define the fitness function as $\Phi(\pi):=C_{\max}+\mu T$, where $T$ is the total tardiness of $\pi$, with $\mu:=n\cdot(\max_{j\in J}t_{j}+\max_{i,j\in J}\ell(\tau_{i},\tau_{j}))+1$ for the job set $J$. Note that $\Phi(\pi)$ essentially models a lexicographic optimization, where the primary goal is to minimize the total tardiness and the secondary goal is to minimize the makespan.
@@ 6.2 Experiments
We perform experiments to evaluate our approaches for the real-world use case from the company CUNA. The products produced by CUNA are available in eight different color/material combinations, all of which are manufactured on the same injection molding machine. For our experiments, we use a $8\times 8$ setup matrix containing the machines setup times of a real-world injection molding machine in minutes as integers.
@@
For all three starting solutions, we evaluated our four hill climbing strategies (Win, Win+Sw, MW, MW+Sw) on 1)1)1 ) a dataset of instances where feasible schedules exist, and 2)2)2 ) a dataset of instances where no feasible schedule exists. Each instance ran within a time limit of 20 minutes. Furthermore, as a baseline, we ran experiments with the two genetic algorithms GAD and MGA and with the PILS1 algorithm. Since GAD, MGA, and PILS1 are non-deterministic randomized algorithms, we performed 20 independent repetitions with a time limit of 20 minutes each and kept the best found results.
We summarize our main experimental findings as follows: First, all hill climbing algorithms provide good results for instances where feasible schedules are possible. The algorithms were mostly able to provide good results even if the starting solution was not feasible. Overall, the window algorithm appears to be well-suited to handle these instances. Second, on instances where no feasible schedules exist, local search via single swaps often has the biggest influence on the solution quality and provides good results: Especially for the large instances with $n\in\{250,300\}$, all improvements were reached by swaps and no parameterized local search step has been applied within the time limit of 20 minutes. On the smaller instances, the VNS algorithms Win+Sw and MW+Sw help the local search via swaps to escape bad swap-optimal solutions and provide good results for this case. Overall, Win+Sw appears to be well-suited to handle these instances. Third, the hill climbing algorithms outperform the baseline genetic algorithms in the solution quality. On the small instances ($n=50$), PILS1 performed slightly better than the hill-climbing algorithms, while on larger instances ($n\geq 100$), the parameterized local search algorithms mostly provide better results than PILS1. This might indicate that parameterized local search is a reasonable alternative to the perturbation of schedules to escape poor local swap-optimal solutions.